# Quantum enhanced Monte Carlo simulation for photon interaction cross sections

### v1

@@ I Introduction
High-energy physics (HEP) aims to uncover the fundamental constituents of matter and their interactions, necessitating increasingly sophisticated computational methods to simulate particle interactions and analyze vast data sets. Large-scale Monte Carlo (MC) simulations are indispensable for connecting observational data with theoretical predictions and for interpreting diverse physical phenomena. Well-known event generators, such as Geant4 utilize Markov chains to model processes that include scattering, decay, and absorption. By accumulating random transitions at each step, they provide accurate predictions of particle spectra, interaction cross sections, and energy budgets. However, these simulations often require extremely high computational costs; achieving adequate precision can require generating millions to billions of random events. As experimental conditions become more complex and collision event counts grow exponentially, the required computational resources and run times escalate dramatically, creating serious bottlenecks in both research and practical applications. Various studies have been conducted to overcome these bottlenecks by integrating parallel computing, deep learning, and advanced sampling techniques [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
In this context, quantum computing, which leverages quantum-mechanical principles to address classically challenging problems, has drawn considerable attention. Through superposition and entanglement, quantum computers promise more efficient exploration of large state spaces and offer the potential for exponential or polynomial speedups in specific tasks. In fact, various quantum algorithms have already been investigated for lattice gauge theories, parton distribution functions, and quantum field simulations in nuclear and high-energy physics, suggesting that the synergy between quantum computing and HEP could spark innovative computational approaches [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Furthermore, it is important to note that this study should be viewed not only as an estimation of photon interaction cross sections but also as a proof-of-concept for simulating photon transport characteristics. In doing so, the proposed quantum algorithm addresses the computational bottlenecks inherent in classical HEP simulations and lays the groundwork for future extensions to more complex physical phenomena, including multiple interaction channels and partial energy emission processes.
Among these, Quantum Amplitude Estimation (QAE) [29] has gained particular interest due to its potential to reduce the complexity and achieve a quadratic speedup over classical Monte Carlo (MC) method. However, a core difficulty lies in converting classical probability distributions into quantum states, a process that can pose significant challenges.
In this paper, we propose an integrated framework that combines a quantum walk (QW) [30, 31] with the QAE algorithm and apply it to a canonical HEP scenario involving high-energy photons interacting with electrons in a water medium. Photon interactions in water provide a suitable demonstration for implementing physical processes on limited quantum hardware and comparing quantum-derived results with classical computation. Quantum walks utilize coin and shift operations to unitarily manage probabilistic branching, enabling the construction of quantum states that encode complex probability distributions. Our framework (1) employs a discrete-time QW to naturally embed the probability distribution at each step into a quantum state, and (2) couples this with amplitude estimation to tackle Markov chain based problems such as interaction cross sections achieving a quadratic speedup. Preserving this stepwise stochastic transition structure implies that the same framework can be broadly extended to problems involving photons, electrons, heavy ions, or even complex probabilistic transport models, including those in fields like finance [32, 33, 34, 35, 36, 37].
This approach directly targets computational bottlenecks in HEP simulations. As noted, Monte Carlo methods remain a cornerstone of high-energy and particle physics, yet their substantial demands on CPU time and memory can delay research or impose constraints. Here, we explore the potential for augmenting or accelerating classical HEP calculations through quantum algorithms. By comparing our proposed quantum algorithm with a standard MC baseline, we show that it correctly reproduces the underlying physics while significantly reducing the effective number of samples (or iterations) needed for a given accuracy. Such findings bridge quantum computing and HEP, indicating that, as quantum devices scale in qubit number and fidelity, one could potentially tackle large-scale particle physics simulations that are currently considered infeasible.
@@
This paper is organized as follows. In Sec. II, we briefly review the theoretical background of quantum walks (QW) and quantum amplitude estimation (QAE). In Sec.,III, we describe the proposed quantum simulation algorithm based on quantum walk and quantum amplitude estimation for photon interactions. In Sec. IV, we present quantum enhanced simulation results using Qiskit, including the speedup achieved compared to classical computing. Finally, Sec.  V summarizes the conclusions of this work and addresses future research directions and potential extensions.
@@ II.2 Quantum Amplitude Estimation
The standard form of QAE is derived from Quantum Phase Estimation (QPE). In Fig. 1(a), QPE circuit with $m$ ancillary qubits applies $Q^{k}$ in an exponentially increasing powers to estimate the amplitude $a$ as
@@ III.1 Theoretical outline of the particle interaction algorithm
where $\rho$ is the material density. For a 10 MeV photon incident on water, contributions other than Compton scattering and pair production amount to less than 1%, making them negligible [44]. Since Compton scattering alone accounts for over 75% of the total attenuation, the subsequent calculations focus exclusively on Compton scattering. Hence,
@@
This can decrease the circuit width by a factor of $\log_{2}(n)$ compared to a model with $n$ separate probabilistic branches. Since fully implementing every branch on IBM’s 32-qubit quantum simulator is not yet feasible, we employ this simplified toy model. This approach is sufficient for proof-of-concept purposes, and once hardware capabilities improve, one can extend the number of possible interaction outcomes per branch from two to an arbitrary integer $k$, enabling more general simulations.
@@ IV Simulation Results
V Conclusion
We propose a novel quantum framework that combines discrete-time quantum walks (QW) and iterative quantum amplitude estimation (IQAE) to model high-energy photon transport in water. By embedding the probabilistic transitions of photons into superposition states, our approach theoretically guarantees a quadratic speedup over classical Monte Carlo (MC) methods while enabling precise amplitude extraction. Benchmark comparisons with classical simulations show that the quantum scheme aligns well with classical models, as evidenced by low mean squared error (MSE) and Kullback–Leibler (KL) divergence values. This confirms the viability of our quantum-walk-based approach in high-energy physics (HEP) simulations and highlights its potential to overcome the computational bottlenecks inherent to large-scale MC techniques.
@@ V Conclusion
However, in this study, we simplify the interaction of 10 MeV photons to a single channel, thereby excluding certain physical details from our model. While this simplification takes into account current hardware constraints, it is important to recognize that real high-energy physics phenomena may involve multiple interaction channels occurring concurrently. In order to capture more realistic physics beyond the single dominant interaction channel, we need to consider multiple branching models that account for additional processes such as pair production and photonuclear reactions. This can be realized by expanding the dimension of the coin used in the quantum walk, thereby allowing multiple interaction channels to be represented within a single step.
Moreover, real high-energy photons are often only partially absorbed in each interaction, losing some fraction of their energy but continuing on with lower energy. Hence, one may split the photon energy into discrete levels in the simulation, extending the standard quantum walk formulation to encode additional energy states. This, however, increases both the number of qubits and the circuit depth.
@@
It is also crucial to incorporate secondary particle generation [50, 51]. This can be done by defining one of the probabilistic branches as secondary particle creation and assigning an ancilla qubit to record the quantum state associated with that event. In doing so, the position, energy, and other parameters of the secondary particle can be tracked. In a hybrid quantum-classical manner, the newly generated secondary particle can be simulated independently as a new track and added to a stack for more realistic simulations. Further research will be required to address limitations related to circuit depth, qubit counts, and quantum noise to realize this approach on real quantum hardware.
Despite these challenges, our results illustrate the promise of quantum-enhanced algorithms in the realm of high-energy physics simulations. By naturally incorporating the Markov-chain-based particle transport process through quantum walks, and combining it with advanced amplitude estimation techniques, this approach provides a scalable pathway to address large-scale HEP simulation problems.
As quantum devices advance in terms of qubit counts, coherence times, and gate fidelities, the methodology presented here is expected to generalize to more complex physical processes. In addition to photon transport, the same framework can be applied to simulate various particles and interactions in nuclear and particle physics, offering a powerful new tool for accelerating Monte Carlo-based analyses across numerous scientific and industrial domains where computational costs are high.
VI Acknowledgements
@@
VII DATA AVAILABILITY
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.

### v2

# Quantum walk based Monte Carlo simulation for photon interaction cross sections

@@ I Introduction
High-energy physics (HEP) seeks to uncover the fundamental constituents of matter and their interactions. However, simulating the intricate processes underlying particle interactions remains a formidable challenge. Classical Monte Carlo (MC) methods, which underpin many simulation tools like Geant4, rely on generating millions to billions of random events to accurately predict particle spectra, interaction cross sections, and energy budgets. This enormous computational burden not only slows down research progress, but also restricts the exploration of complex experimental conditions, posing a serious constraint on advancing HEP studies.
To alleviate these computational bottlenecks, numerous approaches have been proposed, including parallel computing [1, 2, 3], machine learning [4, 5, 6, 7, 8, 9], and advanced sampling techniques [10, 11]. While these methods offer improvements, they are fundamentally limited by the intrinsic challenges of classical simulations. Despite substantial research efforts, classical methods still have not completely solved the problem of computational bottlenecks.
In this context, quantum computing has attracted significant attention because of its potential to solve classically challenging problems. By leveraging quantum entanglement and quantum superposition, quantum computing offers a fundamentally different approach to simulation by enabling a much more efficient exploration of vast state spaces. In fact, research on quantum algorithms has advanced in various areas of high-energy and nuclear physics such as medium effects in dense matter [12, 13], parton showers [14, 15, 16, 17], scattering processes [18, 19, 20], quantum field simulations [21], event analysis [22, 23, 24], jet evolution [25] and dark sector studies [26]. These investigations suggest that the synergy between quantum computing and high-energy physics may lead to innovative computational methods [27, 28]. Moreover, the versatility of these quantum techniques indicates potential applications in other fields, including finance [29, 30, 31, 32, 33, 34, 35], where complex probabilistic transport models are also of great interest.
Among several quantum algorithms, Quantum Amplitude Estimation (QAE) [36] holds the potential to deliver a quadratic speedup over traditional Monte Carlo methods. To apply this approach in practice, classical probability distributions must be mapped to quantum states. This mapping is a significant challenge that impedes the practical implementation of the technique.
In this paper, we propose an integrated framework that combines a discrete time quantum walk [37, 38] with the quantum amplitude estimation algorithm to simulate high-energy photon interactions in water. By embedding the probabilistic structure of particle transport directly into quantum superposition states via the quantum walk, this approach provides a more physically meaningful representation compared to previous overly simplified models [39, 40] or those based on quantum generative adversarial networks [41, 42, 43] and reduces the number of samples required to achieve a given level of accuracy.
In our study, we implemented a simplified model for the interaction of 10 MeV photons in water using the IBM Qiskit quantum simulator. The results indicate that our simulation produces very low mean squared error and Kullback-Leiber divergence values compared with classical Monte Carlo methods. This confirms that the combination of the quantum walk and iterative quantum amplitude estimation provides high accuracy and stability in practice. In particular, our implementation of iterative quantum amplitude estimation reproduces the theoretically guaranteed quadratic speedup and exhibits rapid convergence in experiments. In addition the extension of the quantum walk based interaction model demonstrates consistency with classical simulations for large high-energy physics problems and thus proves the method has robust scalability. Our study can be regarded as a proof of concept for simulating general particle transport properties beyond photon cross section estimation.
The current model assumes a single interaction channel with complete photon absorption, reflecting current hardware limitations. In the future, it will be necessary to incorporate multiple interaction channels, partial energy loss and secondary particle generation to fully capture the complexity of realistic high-energy processes.
@@
This paper is organized as follows. In Sec. II, we briefly review the theoretical background of quantum walks (QW) and quantum amplitude estimation (QAE). In Sec.,III, we describe the proposed quantum simulation algorithm based on quantum walk and quantum amplitude estimation for photon interactions. In Sec. IV, we present simulation results using Qiskit and demonstrates the speedup achieved relative to classical methods. In Sec. V we analyze the results and discuss comparisons with existing approaches, limitations and potential directions for future research. Finally, Sec. VI summarizes the overall achievement of the paper.
@@ II.2 Quantum Amplitude Estimation
The standard form of QAE [36] is derived from Quantum Phase Estimation (QPE). In Fig. 1(a), QPE circuit with $m$ ancillary qubits applies $Q^{k}$ in an exponentially increasing powers to estimate the amplitude $a$ as
@@ III.1 Theoretical outline of the particle interaction algorithm
where $\rho$ is the material density. For a 10 MeV photon incident on water, contributions other than Compton scattering and pair production amount to less than 1%, making them negligible [50]. Since Compton scattering alone accounts for over 75% of the total attenuation, the subsequent calculations focus exclusively on Compton scattering. This simplification was essential for designing circuits that can run on quantum simulators with qubit limitations, representing a deliberate trade-off between physical realism and the constraints of quantum simulator. Hence,
@@
This can decrease the circuit width by a factor of $\log_{2}(n)$ compared to a model with $n$ separate probabilistic branches.Since fully implementing every branch on IBM’s 32-qubit quantum simulator is not yet feasible due to the limited qubit count, we employ this simplified toy model. This approach is sufficient for proof-of-concept purposes, and once hardware capabilities improve, one can extend the number of possible interaction outcomes per branch from two to an arbitrary integer $k$, enabling more general simulations.
@@ IV Simulation Results
V Discusssion
In this study, we implemented a simplified model for the interaction of 10 MeV photons in water media using the IBM Qiskit quantum simulator. Compared to classical Monte Carlo methods, our simulation achieves very low mean squared error (MSE) and Kullback–Leibler (KL) divergence values, confirming that the combination of quantum walks and iterative quantum amplitude estimation (IQAE) provides high accuracy and stability in practice. In particular, amplitude estimation via IQAE demonstrates a theoretical quadratic speedup in our experiments.
In comparison to other recent studies, our work contributes in two main aspects. First, most existing research applying QAE often relies on overly simplified models [39, 40] or quantum generative adversarial network (QGAN)-based methods [41, 42, 43] that do not adequately capture the underlying physical processes. In contrast, our study is the first to introduce a quantum walk (QW) algorithm into the quantum embedding of probability distributions, leading to a more sophisticated and physically meaningful model and experimentally verifying the theoretical quadratic speedup in quantum amplitude estimation. Second, by implementing and scaling up a QW-based interaction model, we confirmed consistent results compared to classical simulations, thereby demonstrating the robust scalability of our method for large-scale high-energy physics (HEP) simulations.
@@ V Discusssion
However, the current model employs a simplified approach by assuming only a single interaction channel with complete photon absorption, and thus does not fully reflect complex phenomena such as multiple interactions, partial energy loss, and secondary particle generation in real high-energy physics scenarios. Although this simplification accounts for current hardware constraints, it is important to recognize that realistic high-energy physics requires the inclusion of these more complex interactions. To capture physics beyond the dominant single interaction channel, multiple branching models incorporating additional processes need to be introduced. This can be achieved by expanding the dimension of the coin in the quantum walk, allowing multiple interaction channels within a single step.
Additionally, high-energy photons frequently lose only a fraction of their energy during each interaction and continue to propagate with reduced energy. Consequently, one can discretize the photon energy into multiple levels in the simulation, extending the standard quantum walk formulation to encode additional energy states. However, this approach increases both the number of qubits and the circuit depth. To implement a physical model with probabilistic branches $k$ per step in a quantum circuit, it is necessary to allocate a number of qubits in both the coin register and the walker register that is $\log_{2}(k)$ times greater than the original requirement. Consequently, incorporating these additional features, given the limitations of current quantum simulator, remains highly challenging.
@@
Incorporating secondary particle generation is also crucial [56, 57]. One probabilistic branch in the simulation can be designated for secondary particle creation, and an ancilla qubit can be assigned to record the corresponding quantum state. This allows the position, energy, and other parameters of the secondary particle to be tracked. A hybrid quantum-classical scheme enables newly created secondary particles to be simulated independently as new tracks, which can be added to a stack for more realistic simulations.
Since our simulations were conducted under ideal conditions using a quantum simulator, practical constraints such as limited qubit counts, circuit depth, and quantum noise on real quantum hardware were not considered. In particular, our quantum circuit involves multiple multi-qubit gates, which pose challenges due to error accumulation when executed on current quantum hardware [58, 59]. As a proof-of-concept, we believe that exploring such approaches is essential in anticipation of future advancements in quantum computing. Despite these challenges, our findings demonstrate that the natural integration of Markov-chain-based particle transport processes via quantum walks, combined with advanced amplitude estimation techniques, offers a scalable approach to address large-scale HEP simulation problems.
VI Conclusion
In this paper, we propose a new quantum framework that combines discrete-time quantum walk (QW) and iterative quantum amplitude estimation (IQAE) to model high-energy photon transport in water. By using the quantum walk algorithm to embed the discrete stochastic process of photon interactions into quantum superposition states and employing the QAE algorithm to extract the statistical features of this process, we confirm that this approach theoretically guarantees a quadratic speedup over classical Monte Carlo (MC) methods while also enabling highly precise amplitude extraction. Quantum simulation results reveal low mean squared error (MSE) and Kullback–Leibler (KL) divergence, reproducing probability distributions nearly identical to those of classical models. This demonstrates that our framework is both valid and feasible in the field of high-energy physics (HEP).
Although the current model has been simplified to accommodate hardware limitations by adopting a single interaction channel and complete absorption, it provides a robust foundation for incorporating more complex interaction channels and multiple energy levels in the future. As quantum devices advance in terms of qubit counts, coherence time, and gate fidelity, the framework presented in this study is expected to extend to handle increasingly complex physical processes. In addition to photon transport, the same method can be applied to diverse particles and interactions in nuclear and particle physics, offering a powerful tool for accelerating Monte Carlo based analyses in scientific and industrial fields where computational costs are high. Ultimately, this work indicates the significant potential of quantum enhanced algorithms in high-energy physics simulations and opens the door to a scalable solution for overcoming the computational bottlenecks inherent in large-scale MC techniques.
Acknowledgements.
@@
Data Availability
The data that support the findings of this article are not publicly available upon publication because the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.